Circular No. 5958

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 1915+105 E. Gérard, Observatoire de Paris-Meudon; L. F. Rodríguez, National

Radio Astronomy Observatory and Universidad Nacional Autónoma de México; and I. F. Mirabel, Service d’Astrophysique, Centre d’Etudes de Saclay, report: “A new radio outburst was detected from the hard xtransient in Aquila on Mar. 24 and 25. The source was observed with the Nançay Radiotelescope and the Very Large Array. The Nançay radio flux was 1.5 and 1.4 Jy at 21 cm and 0.8 and 0.6 Jy at 9 cm, respectively, around Mar. 24.28 and 25.28 UT. Using the VLA around Mar. 24.58, fluxes of 1.10 Jy at 6 cm, 0.81 Jy at 3.6 cm, and 0.55 Jy at 2 cm were measured. This outburst is comparable in intensity to that recorded by Rodríguez and Mirabel in early Dec. 1993 (IAUC 5900). No evidence for linear polarization is found in the radio emission at 21 and 9 cm. The source has an infrared counterpart (Mirabel et al. 1994, A.Ap. 282, L17), and multiwavelength observations are strongly encouraged to follow up the outburst and help understand the nature of GRS 1915+105”.

SUPERNOVAE 1994F, 1994G, 1994H Corrigenda to IAUC 5956: line 6, add H. Pack and M. Cronin, Mt.

Herman School, Northfield, MA; line 13, for R = 22.0 read R = 22; line 14, for 1′′.1 west read 2′′.7 west; line 17, for 1993 Dec. 20 read 1993 Dec. 21; line 18, for J. B. Oke read J. B. Oke, J. Cohen, and T. Bida; line 28, for M. Turrato read M. Turatto; line 30, for redshift of ∼ 0.32 read redshift of 0.373 [the cluster being Abell 0370].

SUPERNOVA 1994D IN NGC 4526 H. Mikuž, Ljubljana, Slovenia, reports the following (comparison stars

from Kilkenny and Malcolm 1984, MNRAS 209, 169): Mar. 19.15 UT, V = 11.86 ± 0.02, B − V = +0.12 ± 0.03, V − R = +0.13 ± 0.03, V − I = −0.10 ± 0.04.

NOVA AQUILAE 1993 Further photometry by Mikuž (cf. IAUC 5794, 5802, 5907): Mar. 19.17

UT, V = 13.43 ± 0.02, B − V = +0.90 ± 0.03.

1994 March 25 Daniel W. E. Green